MONT 104N – Modeling the Environment Paper Assignment 1 – Montaigne and the First Year Summer Reading due: September 14, 2012

As you all know, the book How To Live, or A Life of Montaigne by Sarah Bakewell was chosen this year as the summer reading for the incoming class of 2016 at Holy Cross. Professor Francisco Gago-Jover, your class dean, explained this choice to the faculty in an email message sent out over the summer as follows:

Montaigne (1533-1592), one of the most influential writers of the French Renaissance, is also the creator of the essay as a literary genre. He became famous for his effortless ability to merge in his writings serious intellectual speculation with casual anecdotes and autobiography. His massive volume Essais (literally, “Attempts”) contains, to this day, some of the most widely influential essays ever written. Admired by some later thinkers, reviled by others, Montaigne is remarkably modern even to readers today. His attempt to examine the world through the lens of the only things he can depend on implicitly–his own judgment and personal experience–makes him more accessible to modern readers than any other author of the Renaissance.

The book How to Live: or a Life of Montaigne is a hybrid biography which compares and contrasts biographical details with excerpts from the Essays, thus describing Montaigne’s philosophy using stories from his life. The book also explores different approaches to life derived from Montaigne’s Essays, such as “Be ordinary and imperfect,” “See the world,” “Guard your humanity,” “Wake from the sleep of habit,” “Let life be its own answer,” and perhaps most characteristically for Montaigne, “Question everything.” If there is a central message, it is about individuals trying to find their way through life.

I’m not sure what information you received, but I imagine Professor Gago-Jover said some similar things in his communications with you over the summer. For your first larger-scale writing assignment for this course, you will write an “opinion piece” in which you address the following main question and follow-ups:

Main Question: Did you think this was a successful choice as a summer reading assignment? Were you stimulated to learn (more) about Montaigne and read the Bakewell book by the information you received?

Based on your answer to the Main Question, chose one of the following:

1. If so, did you find what Bakewell had to say about Montaigne interesting and valuable? Which, if any, of the proposed answers to the question “how to live?” did you find to be especially personally compelling or meaningful? Have you made any choices during your time at Holy Cross so far based on those proposed answers? Also, did reading “On Cannibals” add to your understanding of Montaigne (that is over and above what you read in Bakewell)?

2. If not, what was the main issue for you that kept you from seeking out and reading the Bakewell book? Have you already decided you know how to live, for instance? Or did the book just seem too remote and outside of your own experience to draw your interest? Did reading “On Cannibals” make you reconsider your earlier decision, or did it convince you you made the right choice? Most importantly, can you suggest a book or other text that might have been a better choice as a summer reading for the incoming first year class? Or is the whole idea of a summer reading for the first year class not worth pursuing at all?

Please be honest in your answers to these questions. Use correct standard English and a reasonably formal style. Don’t just say what you think I might want to hear, because I don’t have any preconceived ideas about this and the point of this assignment is for you to figure out what you really think and then to express that in a coherent and convincing way.

I will be using this assignment to determine how much we need to work on technical writing issues (grammar, spelling, sentence structure, paragraphing), as well as seeing how effectively you get your points across in writing. So the grading will have two components, one for “form” (the technical issues), and one for “content” (based on the independent thought your essay shows and how well you conveyed that to your reader). If you find the templates from “They Say, I Say” useful as a way to think about structuring the essay, by all means incorporate them.

Important Note: It is perfectly OK with me if you want to share a rough draft of your essay with other students in the class, or with me, to get comments and suggestions. “Testing out” your ideas that way is often an important part of getting them down on paper in an effective form.

Prepare your essay using Microsoft Word and submit as a .doc file attachment to this email address:

                           jlittle@holycross.edu

Your assignment should arrive no later than 5:00pm on Friday, September 14. Aim for somewhere between 3 and 5 double-spaced pages.

For this assignment, your are not expected or required to refer to any sources other than (possibly) Bakewell’s book and the essay we read. If you do want to use direct quotations from the Bakewell book or “Of Cannibals,” identify the source with a notation of the form: (Bakewell, p. 133). Consult with Professor Little if you want to make use of other sources.